OpenStudy (anonymous):

Solve: x2 - 20x + 36 = 0 A. {6, -6} B. {2, 18} C. {-4, -9} D. no solution

OpenStudy (callisto):

you can use quadratic formula to solve it \[x=(-b \pm \sqrt{b^2 - 4ac}) /(2a)\]

OpenStudy (apoorvk):

B

OpenStudy (callisto):

or you can solve it by factoring the expression on the left hand side

OpenStudy (arnavguddu):

x2-20x+36 use a method----- x = (-b+- sqrt(b^2- 4ac))/(2a) a=1 b=-20 c=36 so sqrt((-20)2 - 4(1)(36)) = 16 so x=(20+-16)/2=(36)/2 or (4)/2 = (18, 2) Ans-B

OpenStudy (arnavguddu):

Alternatively---- x^2 - 18x - 2x + 36 = x(x-18) - 2(x-18) = (x-2)(x-18)=0 So either x-2=0 or x-18=0 Thus, x=2 or x=18 Ans x=(2, 18) B